OpenStudy (anonymous):

the pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is 705 feet, and its height is 471 feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work.

OpenStudy (anonymous):

note its a square pyramid not rectangular pyramid

Directrix (directrix):

When I did this before, did I get an incorrect answer? I thought I cranked it out with a square base.

OpenStudy (anonymous):

the square pyramid has a different formula then the rectangular pyramid

OpenStudy (anonymous):

@phi help

Directrix (directrix):

Now, what is wrong with this work? Volume = (1/3) Bh where B is the area of the base and h is the height V = 1/3 (705)^2 (471) V = 78, 032, 925 cubic feet. Note the (705)^2 which is the area of the square base. http://openstudy.com/users/cool987654321#/updates/4f70825de4b0eb858773121d

OpenStudy (anonymous):

someone sent me message and said its wrong and its for rectangular pyramids, but if you are sure of ur answer ill go with it thx

OpenStudy (phi):

type (705)^2 (471)/3= in google search window

Directrix (directrix):

It's not about answers for me, it's about technique and your understanding that the B in the formula is a place-holder for whatever polygon is on the base of the pyramid. Note that in this problem I worked for you, the B for base slot reflects that the formula for the area of a non-square rectangle was used in the B slot. ( Note: All squares are rectangles so that's why I wrote non-square rectangle.) V = (1/3) B h where B is area of the Base and h is the height of the pyramid. 756 = (1/3) 9 (14) h --> 9*14 is area of rectangular base <--- length times width here 756 = 3 *14 * h 756/(3*14) = h 756/42 = h 18 = h ----> Check my work. http://openstudy.com/users/directrix#/updates/4f708846e4b0eb85877319ea

Directrix (directrix):

Note that I have laid out my work for anyone to help find errors. Perhaps your "math whisperer" will do the same. :)